OpenStudy (anonymous):

change 144% as a mixed number

OpenStudy (anonymous):

change 144% as a mixed number and reduce if possible

OpenStudy (anonymous):

1 and 11/25

OpenStudy (anonymous):

1.44 is the decimal number you'll have to write 0.44 as a fraction

OpenStudy (anonymous):

0.44 = 44/100 44/4 =11 100/4 = 25 11/25 is the smallest way to put it because 11 is prime

OpenStudy (anonymous):

thanks yall

OpenStudy (anonymous):

change 3-1/2 as a fraction and reduce if possible